# Advancing Mobile GUI Agents: A Verifier-Driven Approach to Practical Deployment

### v2

@@ Abstract.
V-Droid sets a new state-of-the-art task success rate across several public mobile task automation benchmarks: 59.5% on AndroidWorld, 38.3% on AndroidLab, and 49% on MobileAgentBench, surpassing existing agents by 9.5%, 2.1%, and 9%, respectively. Furthermore, V-Droid achieves an impressively low latency of 0.7 seconds per step, making it the first mobile agent capable of delivering near-real-time, effective decision-making capabilities.
@@ 1. Introduction
Figure 1. Task success rate and decision making latency per step of current SOTA mobile agents   [^2] The task success rate achieved by GPT-4 and 4o, Claude, Llama, Gemini and Qwen is measured with the default prompt templates i.e., T3A and M3A, in AndroidWorld benchmark.and V-Droid evaluated on AndroidWorld benchmark. The latency of 2B, 7B and 8B agents are measured on 2×2$\times$2 × Nvidia 4090. For 72B or MoE agents, the latency is measure on 4×4$\times$4 × Nvidia A100 80G.
Controlling mobile devices via natural language has been a longstanding aspiration for researchers and developers in the mobile domain (Wen et al., 2024a; Lee et al., 2024; Wen et al., 2024b; Shvo et al., 2021), promising opportunities to automate repetitive tasks and elevate user convenience. Unlike API-based agents that rely on predefined function calls (Li et al., 2023; Chen and Li, 2024), mobile GUI agents simulate human interactions, allowing them to operate across diverse applications through Graphical User Interface (GUI). However, developing such an agent poses significant challenges: it needs to not only interpret on-screen content but also make reasonable decisions to execute multi-step tasks within dynamic and complex GUI environments.
In recent years, a variety of LLM-powered mobile GUI agents have been proposed (Wen et al., 2024a; Agashe et al., 2025; Qin et al., 2025; Gou et al., 2025; Zheng et al., 2024; Wang et al., 2024c, e; Hong et al., 2024; Yang et al., 2024). These agents typically utilize LLMs as generators, generating decisions (e.g., reasoning and actions) based on the current task states (e.g., user interfaces, task descriptions), leveraging the contextual understanding and reasoning abilities inherent to LLMs. These agents, however, fall short of meeting practical deployments. Fig.1 presents the performance of state-of-the-art (SOTA) mobile agents in terms of task success rate (SR) and decision-making latency per step, evaluated on the AndroidWorld benchmark(Rawles et al., 2024). As shown, the highest SR achieved by existing agents is only 50%, significantly lower than the human performance of 80%. Additionally, latency remains a critical challenge. Agent-S2 (Agashe et al., 2025), powered by GPT-4o and UI-TARS-72B (Qin et al., 2025), takes nearly 15 seconds on average to make a single action decision.
@@
The fundamental research problem in developing practical mobile agents lies in enhancing their decision-making capabilities for mobile tasks while simultaneously maintain reasonable decision-making latency. To achieve this goal, we introduce V-Droid, which, as illustrated in Fig. 1, achieves a marked improvement in task success rate, reaching a new SOTA at 59.5% on the AndroidWorld benchmark, while significantly reducing the decision making time to less than a second.
The key idea behind V-Droid is transforming the paradigm of mobile agents by using LLM as verifiers instead of generators, which is illustrated in Fig. 2: rather than directly generating the final decision, when LLMs are used as verifiers, potential actions are first extracted. The verifier-driven agents then explicitly evaluate each candidate action e.g., by prompting, ”Is action X helpful for completing the task?” Following these evaluations, the agent selects the action with the highest estimated value, e.g., the likelihood of generating a ’Yes’ token.
@@
Pair-wise process preference training method. To improve the decision-making capabilities of V-Droid, we introduce a pair-wise preference training strategy tailored for verifiers. This method enables the verifier to prioritize correct actions by assigning them higher scores while penalizing incorrect actions at each task step. Furthermore, unlike prior post-training approaches that rely solely on mobile GUI data, our training leverages labeled mobile task trajectories with fine-grained process supervision (Lightman et al., 2023). This approach significantly enhances the task-specific decision-making proficiency of V-Droid.
@@
The verifier in V-Droid, which uses the small language model (SLM), Llama-3.1-8B as the backbone, is trained using the $P^{3}$ method across four iterative training rounds with 110K samples collaboratively annotated by humans and agents. We evaluate V-Droid on three public, realistic task benchmarks: AndroidWorld (Rawles et al., 2024), AndroidLab (Xu et al., 2024), and the MobileAgentBench (Wang et al., 2024a). On these benchmarks, V-Droid sets new SOTA task success rates of 59.5%, 38.3%, and 49%, surpassing previous best-performing agents by absolute margins of 9.5%, 2.1%, and 9.0%, respectively. Compared to other SLM-based agents, V-Droid achieves significant SR improvements of 50%, 22%, and 15% on these benchmarks. In addition, V-Droid consistently delivers a 10–30$\times$ speedup over prior SOTA mobile agents. Notably, using only two Nvidia 4090 GPUs, V-Droid achieves a decision-making latency of just 0.7 second, making V-Droid the first mobile agent with near-real-time response capability.
In summary, we make the following contributions in this paper:
@@
  • V-Droid significantly outperforms previous SOTA in task success rates on multiple public benchmarks while reducing latency by 10–30$\times$, achieving the first near-real-time response for mobile agents.
@@ 2.1. LLM Powered Mobile GUI Agent
Figure 2. The key differences in agent architecture between using LLMs as generators and using LLMs as verifiers in decision-making process: rather than directly determine actions based on task states, verifier-driven agents explicitly evaluate each action before arriving at the final decision.
Recently, the emergence of numerous LLM-powered mobile agents (Wen et al., 2024a; Wang et al., 2024d; Ding, 2024; Wang et al., 2024b; Li et al., 2024b, b) has been observed. The introduction of large language models (LLMs) has significantly improved the ability of mobile agents to comprehend context and generate effective action sequences. In particular, existing agents typically employ the following strategies to enhance performance.
Most mobile agents (Zheng et al., 2024; Wen et al., 2024a; Wang et al., 2024d; Ding, 2024; Wang et al., 2024b; Li et al., 2024b, b) leverage prompt engineering techniques to optimize task execution. For instance, SeeAct (Zheng et al., 2024) employs a ReAct-style (Yao et al., 2023) prompt to break down tasks into manageable steps, thereby reducing errors and mitigating hallucinations in the LLM’s outputs. Additionally, agents like AutoDroid (Wen et al., 2024a) and MobileGPT (Lee et al., 2024) collect task completion traces during offline preprocessing stages for specific applications. These traces are integrated with the memory of pre-trained LLMs, enabling enhanced performance tailored to particular applications. Moreover, some agents utilize a two-stage approach, where advanced grounding models such as UGround (Gou et al., 2025), Aria-UI (Yang et al., 2024), UI-TARS (Qin et al., 2025), OmniParser (Wan et al., 2024; Yu et al., 2025), and Ferret-UI (You et al., 2024), etc., are first employed to comprehensively interpret the UI. This is followed by leveraging pre-trained LLMs for reasoning and action execution. These methods primarily focus on harnessing the inherent capabilities of LLMs for mobile tasks. However, despite the rapid progress in LLMs driven by extensive pre-training, the absence of the training corpus specifically tailored for mobile GUI tasks leaves existing models insufficiently equipped to handle mobile task automation effectively, as demonstrated in Fig. 1.
In addition to utilizing pre-trained LLMs for mobile agents, researchers have explored GUI fine-tuning strategies to improve task execution (Liu et al., 2024; Qwen et al., 2025). This involves adapting a pre-trained language model using domain-specific data, such as annotated screenshots and GUI representations. While GUI fine-tuning enhances the model’s ability to accurately interpret and interact with GUI elements (Zhang et al., 2024b; Gao et al., 2024), it remains insufficient for facilitating task-specific reasoning and multi-step decision-making, especially in dynamic and complex GUI environments, as evidenced in Fig. 1.
@@ 2.2. Opportunity and Challenges
We aim to tackle this challenge by proposing a novel approach: using LLMs as verifiers instead of generators for mobile agents.
@@
Fig. 2 illustrates the key difference in agent architecture between using LLMs as generators and using LLMs as verifiers. The feasibility of verifier-driven agents clearly depends on the presence of an enumerable and extractable action space. In the context of GUI automation tasks on mobile devices, this prerequisite is entirely met. Owing to the constrained screen size and the inherent interaction patterns of touchscreen interfaces, both the types of actions and the number of interactive UI elements on a single page are limited. Fig.3 illustrates the distribution of interactive UI elements within each GUI state during mobile task execution. As depicted, the interactive action space on mobile devices is generally constrained to approximately $20$ actions on average. Although we have a unique opportunity to build verifier-driven agents for mobile tasks, several technical challenges are not well addressed.
Firstly, the detailed architecture and workflow of the verifier-driven agent remain ambiguous. In particular, constructing a well-defined action space is a nontrivial challenge. Solely extracting actions from the user interface is insufficient, as certain useful actions may not be explicitly visible, e.g., navigate home. Furthermore, some actions are inherently continuous rather than discrete, e.g., type text. Beyond that, efficiently verifying a set of actions also poses difficulties, especially considering the best utilizing the hardware parallelism.
Secondly, directly utilizing a pre-trained LLM as the verifier is not a feasible solution for mobile agents. For instance, Llama-3.1-8B, fail to complete any tasks within the benchmark (detailed in § 6.4). Even powerful LLMs, i.e., GPT-4 as the verifier, achieve only $34.5$% task success rate on the AndroidWorld benchmark. The challenge of effectively fine-tuning such a verifier remains unresolved.
@@ 3. Workflow of V-Droid
Figure 4. The Workflow of V-Droid: ① Extracting actions from UI and supplementing default actions; ② Constructing verification prompts with the template for each candidate action; ③ Scoring with the verifier in batch with prefix caching; ④ Completing and executing the selected action; ⑤ Updating the working memory.
Fig. 4 illustrates the workflow of V-Droid. As a verifier-driven agent, V-Droid requires the enumeration of candidate actions prior to estimating the optimal action at each step. To achieve this, V-Droid initially employs a rule-based action candidate extractor from the UI representations. Subsequently, essential action candidates that are not explicitly visible within the current UI (e.g., navigate home) are incorporated. Each action candidate is then encapsulated using a predefined prompt template. Following this, the fine-tuned LLM as the verifier is utilized to evaluate and assign scores to these candidates in batch. The action with the highest score is selected and executed. In the following, we provide an in-depth explanation of each module.
@@ 3.1. Constructing Action Space
A task $\kappa$ is automated through a sequence of steps. At any given step $t$, we aim to extract a set of actions $\mathcal{A}_{\kappa}(t)$, representing the potential interactions available within the UI state $\mathcal{S}_{\kappa}(t)$ associated with the current step $t$, collectively defining the action space. This action space generally comprises two categories: UI-dependent and -independent actions, the latter being default actions that are not explicitly visible.
@@
To identify interactive UI elements, the XML representation of the UI state $\mathcal{S}_{\kappa}(t)$ is directly analyzed, extracted using the Android Accessibility Service (Android Developers, 2025a). A rule-based extractor is applied to detect UI elements that are clickable, long-clickable, scrollable, or editable. These identified elements are then mapped to their corresponding basic action types, thereby forming the UI-dependent action space.
It is worth noting that when a keyboard is displayed on the screen, numerous interactive UI elements may correspond to individual keys, significantly increasing the size of the action space. To address this, V-Droid disables the standard on-screen keyboard and performs text input directly via Android commands (Android Developers, 2025b).
In addition to the UI-dependent actions, there exist actions that are not visible in the current UI but are essential for device control and task completion. Specifically, we supplement every action space $\mathcal{A}_{\kappa}(t)$ with the following default actions: open apps, wait (perform no action but pause for the device response), navigate home (simulate pressing the home button), navigate back (simulate pressing the back button), complete task (indicate that the task has been successfully completed), and answer users’ question (summarize and respond to task-specific queries).
@@ 3.2. Scoring with Verifier
In the role component, we outline the guidelines for operating mobile devices, as in (Rawles et al., 2024). The goal component specifies the task to be automated. The working memory maintains a record of action histories and trajectories for the current task, and it is updated after the execution of each action. For the UI component, we provide a streamlined description of the UI in HTML format, as in  (Wen et al., 2024a). The instruction component delivers explicit directives for the generative verifier, specifying that it must respond strictly with ’Yes’ or ’No’. Finally, the prompt concludes with the specific question: Is the action $\alpha$ helpful for completing the given user task?
@@
Finally, after evaluating all action candidates, the action $\alpha_{\kappa}^{y}(t)$ assigned the highest score is selected for execution at the step $t$:
@@ 3.4. Completing and Executing Action
Before the selected action is exactly executed, three specific types of actions—open app, type text, and answer—require additional completion to specify the target app to open, the content to type, and the response to the user’s query. Thus we employ an LLM to generate the necessary content by prompting the current UI state and working memory.
@@ 4. Pairwise Process Preference Training
The generative verifier in V-Droid, which assigns a score to each action, enables the agent to discern which actions are more likely to contribute effectively to completing the given task. Before presenting our proposed training methods, we first examine shortcomings of existing approaches.
A straightforward approach involves directly employing a pre-trained LLM as the verifier without any post-training, using the logits or the probability of the ’Yes’ token from the output space as the verification score. This method, referred to as LLM-as-a-Judge (Gu et al., 2025), has been evaluated in our experiments (§6.4). However, our measurements indicate that directly leveraging the Llama-3.1-8B as the verifier results in a success rate as low as 0% on the AndroidWorld benchmark (Rawles et al., 2024). Alternatively, researchers have attempted to fine-tune pre-trained LLMs, such as Qwen2.5-VL-72B (Qwen et al., 2025), using Android GUI data (Rawles et al., 2023) through SFT. Unfortunately, the results reveal that the success rate is just 35%.
The rationale behind the suboptimal performance is twofold. First, for the verifier without post-training, we observe that the scores derived from the token space are insufficiently distinguishable, particularly in scenarios with numerous action candidates. Furthermore, simple fine-tuning on GUI data fails to enhance this indistinguishability. Therefore, we believe that preference training is critical to bridging the gap between the generated token space and the desired scoring (reward) space. More importantly, this training must be task-specific rather than relying solely on domain adaptation with GUI data, as the decision-making capabilities of the mobile agent cannot be effectively enhanced otherwise.
We propose the Pairwise Process Preference ($P^{3}$) training method to significantly enhance the decision-making capabilities of verifier-driven mobile agents.

### v3

@@ Abstract.
V-Droid obtains a substantial task success rate across several public mobile task automation benchmarks: 59.5% on AndroidWorld, 38.3% on AndroidLab, and 49% on MobileAgentBench, surpassing existing agents by 5.2%, 2.1%, and 9%, respectively. Furthermore, V-Droid achieves a remarkably low latency of 4.3s per step, which is $6.1\times$ faster compared with existing mobile agents. The source code is available at https://github.com/V-Droid-Agent/V-Droid.
@@ 1. Introduction
Figure 1. Task success rate and latency per step of current mobile agents   [^2] The task success rate achieved by GPT-4 and 4o, Claude, Llama, Gemini and Qwen is measured with the default prompt templates i.e., T3A and M3A, in AndroidWorld benchmark.and V-Droid evaluated on AndroidWorld benchmark. The latency of 2B, 7B and 8B agents are measured on $2\times$ Nvidia 4090. For 72B or MoE agents, the latency is measure on $4\times$ Nvidia A100 80G.
Controlling mobile devices via natural language has been a longstanding aspiration in the mobile domain (Wen et al., 2024a; Lee et al., 2024; Wen et al., 2024b; Shvo et al., 2021), promising opportunities to automate repetitive tasks and elevate user convenience. Unlike API-based agents that rely on predefined function calls (Li et al., 2023; Chen and Li, 2024), mobile GUI agents simulate human interactions, allowing them to operate across diverse applications through Graphical User Interface (GUI). However, developing such an agent poses significant challenges: it needs to not only interpret on-screen content but also make reasonable decisions to execute multi-step tasks within dynamic and complex GUI environments.
In recent years, a variety of LLM-powered mobile GUI agents have been proposed (Wen et al., 2024a; Agashe et al., 2025; Qin et al., 2025; Gou et al., 2025; Zheng et al., 2024; Wang et al., 2024c, e; Hong et al., 2024; Yang et al., 2024). These agents typically utilize LLMs as generators, generating decisions (e.g., reasoning and actions) based on the current task states (e.g., user interfaces, task descriptions), leveraging the contextual understanding and reasoning abilities inherent to LLMs. These agents, however, fall short of meeting practical deployments. Fig.1 presents the performance of state-of-the-art (SOTA) mobile agents in terms of task success rate (SR) and stepwise latency on the AndroidWorld benchmark(Rawles et al., 2024). As shown, the highest SR achieved by existing agents is only 54.3%, significantly lower than the human performance of 80%. Additionally, latency remains a critical challenge. Agent-S2 (Agashe et al., 2025), powered by GPT-4o and UI-TARS-72B (Qin et al., 2025), takes over 25 seconds for a single step.
@@
The fundamental research problem in developing practical mobile agents lies in enhancing their decision-making capabilities for mobile tasks while simultaneously maintain reasonable latency. To achieve this goal, we introduce V-Droid, which, as illustrated in Fig. 1, achieves a marked improvement in task success rate, refreshing the record to 59.5% on the AndroidWorld benchmark, while achieves $6.1\times$ speed up on the step-wise latency.
The key idea behind V-Droid is transforming the paradigm of mobile agents by using LLM as verifiers instead of generators, which is illustrated in Fig. 2: rather than directly generating the final decision, when LLMs are used as verifiers, potential actions are first extracted. The verifier-driven agents then explicitly evaluate each candidate action e.g., by prompting, ”Is action X helpful for completing the task?” Following these evaluations, the agent selects the action with the highest value, e.g., the likelihood of generating a ’Yes’ token.
@@
Figure 2. The key differences in agent architecture between using LLMs as generators and as verifiers for decision-making: rather than directly determine actions based on states, verifier-driven agents explicitly evaluate each action before arriving at the decision.
Pair-wise process preference ($P^{3}$) training method. To improve the decision-making capabilities of V-Droid, we introduce a pair-wise preference training strategy tailored for verifiers. Unlike prior post-training approaches that rely solely on mobile GUI data, our training leverages labeled mobile task trajectories with fine-grained process supervision (Lightman et al., 2023). It enables the verifier to prioritize correct actions by assigning them higher scores while penalizing incorrect actions at each task step. This approach significantly enhances the task-specific decision-making proficiency of V-Droid.
@@
[…]
In summary, we make the following contributions:
@@
[…]
@@ 2.1. LLM Powered Mobile GUI Agent
Recently, the emergence of numerous LLM-powered mobile agents (Wen et al., 2024a; Wang et al., 2024d; Ding, 2024; Wang et al., 2024b; Li et al., 2024b, b) has been observed. The introduction of large language models (LLMs) has significantly improved the ability of mobile agents to comprehend context and generate effective actions. Existing agents typically employ the following strategies to enhance performance.
Most mobile agents (Zheng et al., 2024; Wen et al., 2024a; Wang et al., 2024d; Ding, 2024; Wang et al., 2024b; Li et al., 2024b, b) leverage prompt engineering techniques to optimize task execution. For instance, SeeAct (Zheng et al., 2024) employs a ReAct-style (Yao et al., 2023) prompt to break down tasks into manageable steps, thereby reducing errors and mitigating hallucinations in the LLM’s outputs. Additionally, agents like AutoDroid (Wen et al., 2024a) and MobileGPT (Lee et al., 2024) collect task completion traces during offline preprocessing stages for specific applications. These traces are integrated with the memory of pre-trained LLMs, enabling enhanced performance tailored to particular applications.
Several recent works, such as UGround (Gou et al., 2025), Aria-UI (Yang et al., 2024), UI-TARS (Qin et al., 2025), and Ferret-UI (You et al., 2024), propose training grounding models that leverage pre-trained LLMs to comprehensively interpret and interact with the UI. These approaches exploit the inherent reasoning capabilities of LLMs for mobile tasks. However, despite the rapid advances in LLMs driven by large-scale pre-training, the lack of training corpora specifically tailored to mobile GUI tasks leaves existing models insufficiently prepared for reliable mobile task automation, as demonstrated in Fig. 1.
In addition to utilizing pre-trained LLMs for mobile agents, researchers have explored GUI fine-tuning strategies to improve task execution (Liu et al., 2024; Qwen et al., 2025). This involves adapting a pre-trained language model using domain-specific data, such as annotated screenshots and GUI representations. While GUI fine-tuning enhances the model’s ability to accurately interpret and interact with GUI elements (Zhang et al., 2024; Gao et al., 2024), it remains insufficient for facilitating task-specific reasoning and multi-step decision-making, especially in dynamic and complex GUI environments, as evidenced in Fig. 1.
@@ 2.2. Opportunity and Challenges
We tackle this challenge by proposing a novel approach: using LLMs as verifiers instead of generators for mobile agents.
@@
Fig. 2 illustrates the key difference in agent architecture between using LLMs as generators and using LLMs as verifiers. The feasibility of verifier-driven agents depends on the presence of an enumerable and extractable action space. In the context of GUI automation tasks on mobile devices, this prerequisite is entirely met. Owing to the constrained screen size and the inherent interaction patterns of touchscreen interfaces, both the types of actions and the number of interactive UI elements on a single page are limited.
Fig.3 illustrates the distribution of interactive UI elements within each GUI state. As depicted, the interactive action space on mobile devices is generally constrained to approximately $20$ elements on average. Although we have a unique opportunity to build verifier-driven agents for mobile tasks, several technical challenges are not well addressed.
Firstly, the detailed architecture and workflow of the verifier-driven agent remain ambiguous. In particular, constructing a well-defined action space is a nontrivial challenge. The supported action type varies on different UI elements and certain useful actions may not be explicitly visible, e.g., navigate home. Furthermore, some actions are inherently continuous rather than discrete, e.g., type text. Beyond that, efficiently verifying a set of actions also poses difficulties, especially considering the best utilizing the hardware parallelism.
Secondly, directly utilizing a pre-trained LLM as the verifier is not a feasible solution for mobile agents. For instance, Llama-3.1-8B, fail to complete any tasks within the benchmark (detailed in § 6.4). Even powerful LLMs, i.e., GPT-4 as the verifier, achieve only 34.534.5% task success rate on the AndroidWorld benchmark. The challenge of effectively fine-tuning such a verifier remains unresolved.
@@ 3. Workflow of V-Droid
Figure 4. The Workflow of V-Droid: ① Extracting actions from UI and supplementing default actions; ② Constructing verification prompts with the template for candidate actions; ③ Scoring with the verifier in batch with prefix caching; ④ Completing and executing the selected action; ⑤ Updating the working memory.
Fig. 4 illustrates the workflow of V-Droid. As a verifier-driven agent, V-Droid requires the enumeration of candidate actions prior to estimating the optimal action at each step. To achieve this, V-Droid employs a rule-based action candidate extractor to obtain actions from the current UI state. Each action candidate is then encapsulated using a predefined prompt template. Following this, the fine-tuned LLM as the verifier is utilized to evaluate and assign scores to these candidates in batch. The action with the highest score is selected and executed. In the following, we provide an in-depth explanation of each module.
@@ 3.1. Constructing Action Space
A task $\kappa$ is automated through a sequence of steps. At any given step $t$, we extract a set of actions $\mathcal{A}_{\kappa}(t)$, representing the potential interactions available within the UI state $\mathcal{S}_{\kappa}(t)$ associated with the current step $t$, collectively defining the action space. This action space generally comprises two categories: UI-dependent and -independent actions, the latter being default actions that are not explicitly visible.
@@
To identify interactive UI elements, the XML representation of the UI state $\mathcal{S}_{\kappa}(t)$ is directly analyzed, extracted using the Android Accessibility Service (Android Developers, 2025a). A rule-based extractor is applied to detect UI elements, such as button, checkbox, and textbox, that are clickable, long-clickable, scrollable, or editable. These identified elements are then mapped to their corresponding basic action types. For example, and editable text box $\beta$ would be mapped to the action ‘input {content} to $\beta$. To prevent a combinatorial explosion of candidates, the standard on-screen keyboard is disabled, and text input is instead performed directly via Android commands (Android Developers, 2025b).
In addition to the UI-dependent actions, there exist actions that are not visible in the current UI but are essential for device control and task completion. Specifically, we supplement every action space $\mathcal{A}_{\kappa}(t)$ with the following default actions: open apps, wait, navigate home, navigate back, complete task, and answer users’ question.
The analysis presented in § 7 demonstrates that the rule-based extraction employed by V-Droid effectively encompasses the majority of interactions observed in real-world mobile tasks, thereby ensuring comprehensive practical coverage without unnecessarily complicating the action space.
@@ 3.2. Scoring with Verifier
In the role component, we outline the guidelines for operating mobile devices, as in (Rawles et al., 2024). The goal component specifies the task to be automated. The working memory maintains a record of action histories and trajectories for the current task, and it is updated after the execution of each action. For the UI component, we provide a streamlined description of the UI in XML format, as in  (Wen et al., 2024a). The instruction component delivers explicit directives for the generative verifier, specifying that it must respond strictly with ’Yes’ or ’No’. Finally, the prompt concludes with the specific question: Is the action $\alpha$ helpful for completing the given user task?
@@
Finally, after evaluating action candidates, the action $\alpha_{\kappa}^{y}(t)$ assigned the highest score is selected for execution at $t$:
@@ 3.4. Completing and Executing Action
Before the selected action is exactly executed, three specific types of actions, open app, type text, and answer, require additional completion to specify the target app to open, the content to type, and the response to the user’s query. Thus we employ an LLM to generate the necessary content by prompting the current UI state and working memory.
@@ 4. Pairwise Process Preference Training
The generative verifier in V-Droid, which assigns a score to each action, enables the agent to discern which actions are more likely to contribute effectively to completing the given task. A straightforward approach involves directly employing a pre-trained LLM as the verifier without any post-training, using the logits or the probability of the ’Yes’ token from the output space as the verification score. However, our measurements (§6.4) indicate that directly leveraging the Llama-3.1-8B as the verifier results in a success rate as low as 0% on AndroidWorld benchmark (Rawles et al., 2024).
The rationale behind the suboptimal performance is twofold. First, for the verifier without post-training, we observe that the scores derived from the token space are insufficiently distinguishable, particularly in scenarios with numerous action candidates. Furthermore, simple fine-tuning on GUI data fails to enhance this indistinguishability. For instance, Qwen2.5-VL-72B (Qwen et al., 2025) fine-tuned with Android GUI data (Rawles et al., 2023) achieves only 35% success rate. Therefore, we believe that, to bridge the gap between the generated token space and the desired scoring (reward) space, additional training is necessary. More importantly, this training must be task-specific rather than relying solely on domain adaptation with GUI data, as the decision-making capabilities of mobile agent cannot be effectively enhanced otherwise. We propose the Pairwise Process Preference ($P^{3}$) training to significantly enhance the decision-making capabilities of verifier-driven mobile agents.